# Unsupervised Online 3D Instance Segmentation with Synthetic Sequences and Dynamic Loss

### v1

@@ II-A Unsupervised 3D Instance Segmentation
3D instance segmentation has evolved rapidly over the past few years, with methods broadly falling into proposal-based [3], proposal-free (grouping/clustering), kernel-based [34], and transformer-based paradigms [5, 35]. Proposal-based methods like 3D-SIS first detect object bounding boxes in 3D or via fused RGB-D input, then refine to instance masks, but these often depend strongly on the quality of the 3D detection step [3]. Grouping or proposal-free methods, such as SGPN [36] / PointGroup [37], predict per-point features (embeddings, center offsets, or affinities) and then cluster or group them to form instances. For example, RPGN [38] introduces a notion of regional purity to help mitigate over- and under‐segmentation in clustering. Mask3D [5] is a Transformer-based model for 3D semantic instance segmentation that uses instance queries to directly predict instance masks from point clouds, eliminating the need for hand-crafted voting or grouping mechanisms.
@@ IV-E Ablation Study
TABLE VI: Influence of different sampling strategies for training. “RTO” denotes Reversing the Temporal Order; “NFS” means Non-adjacent Frame Selection.
@@
Effect of Different Sampling Strategies. Table VI evaluates the impact of our proposed sampling strategies during training. Removing either reversing the temporal order (RTO) or non-adjacent frame selection (NFS) consistently lowers performance across all metrics. Without RTO, the scores drop slightly to 0.521 in temporal association, 0.599 in $\mathrm{IoU^{*}}$, and 0.723 in association score. Similarly, excluding NFS reduces the scores further to 0.518, 0.597, and 0.720, respectively. The full model with both strategies enabled achieves the best results, confirming that RTO and NFS are complementary and together improve temporal modeling and overall segmentation quality.
@@
IV-F Qualitative Results
@@
Figure 6: Visual results on PandaSet-GT dataset. Our method produces cleaner and more consistent instance masks compared with pseudo-labels and UNIT. Best viewed in zoom.
@@
Figs 5 and 6 present qualitative comparisons on the nuScenes and PandaSet-GT datasets. Our method yields more accurate and temporally consistent instance segmentation compared with pseudo-labels and UNIT. In nuScenes, UNIT introduces inconsistent predictions (highlighted with red circles), whereas our approach produces temporally consistent masks with clearer object boundaries. On PandaSet, which contains complex urban scenes with diverse object types, our method better preserves object shapes and reduces over-segmentation, leading to results that are closer to the ground-truth annotations. These improvements highlight the robustness and generalization of our approach across different datasets.

### v2

@@ II-A Unsupervised 3D Instance Segmentation
3D instance segmentation has evolved rapidly over the past few years, with methods broadly falling into proposal-based [3], proposal-free (grouping/clustering), kernel-based [34], and transformer-based paradigms [5, 35]. Proposal-based methods like 3D-SIS first detect object bounding boxes in 3D or via fused RGB-D input, then refine to instance masks, but these often depend strongly on the quality of the 3D detection step [3]. Grouping or proposal-free methods, such as SGPN [36] / PointGroup [37], predict per-point features (embeddings, center offsets, or affinities) and then cluster or group them to form instances. For example, RPGN [38] introduces a notion of regional purity to help mitigate over- and under‐segmentation in clustering. Mask3D [5] is a Transformer-based model for 3D semantic instance segmentation that uses instance queries to directly predict instance masks from point clouds, eliminating the need for hand-crafted voting or grouping mechanisms. COSeg [39] improves generalization by refining query-prototype correlations, while GFS-VL [40] combines pseudo-labels from vision-language models with few-shot samples for better novel class segmentation. Seal [41] distills vision foundation models (VFMs) for segmenting diverse automotive point cloud sequences.
@@ IV-E Ablation Study
TABLE VI: Influence of different sampling strategies for training. “RTO” denotes Reversing the Temporal Order; “NFS” means Non-adjacent Frame Selection. “Fixed-gap” denotes non-random fixed-gap sampling where interval is 2.
@@
Fixed-gap 0.508 0.588 0.711
Effect of Different Sampling Strategies. Table VI evaluates the impact of our proposed sampling strategies during training. Removing either reversing the temporal order (RTO) or non-adjacent frame selection (NFS) consistently lowers performance across all metrics. Without RTO, the scores drop slightly to 0.521 in temporal association, 0.599 in $\mathrm{IoU^{*}}$, and 0.723 in association score. Similarly, excluding NFS reduces the scores further to 0.518, 0.597, and 0.720, respectively. The full model with both strategies enabled achieves the best results, confirming that RTO and NFS are complementary and together improve temporal modeling and overall segmentation quality. And our proposed sampling strategy also outperforms the fixed‑gap sampling across all three metrics.
@@
Effectiveness of Point Cloud Sequence Synthesis. As shown in Table VIII, the results highlight that our approach outperforms the Copy-Paste method across all metrics, including $\mathrm{IoU^{*}}$ and $\mathrm{S_{assoc}}$. We observe the highest performance with a substantial improvement in $\mathrm{IoU^{*}}$ (0.589 vs. 0.575) and $\mathrm{S_{assoc}}$ (0.716 vs. 0.708).
TABLE VIII: Comparison between the common copy-paste strategy and our point cloud sequence synthesis method.
Copy-Paste Ours $\mathrm{S_{assoc}^{temp}}$ $\mathrm{IoU^{*}}$ $\mathrm{S_{assoc}}$
$\times$ $\times$ 0.482 0.568 0.696
$\checkmark$ $\times$ 0.496 0.575 0.708
$\times$ $\checkmark$ 0.505 0.589 0.716
TABLE IX: Efficiency and performance comparison on the SemanticKITTI validation split. Model size, inference latency, frames per second (FPS), and $\mathrm{S_{assoc}^{temp}}$ are reported.
Method Latency (s) FPS Training Time (h) Memory (GB) Model Size(M) $\mathrm{S_{assoc}^{temp}}$
TARL-Seg [8] 0.97 1.03 0 0 0 0.231
4D-Seg [6] 1.52 0.65 0 0 0 0.421
3DUIS++ [9] 74.09 0.01 189 0.98 167 0.447
UNIT [6] 0.49 2.01 48 4.16 909 0.482
Ours 0.49 2.01 50 4.16 909 0.523
IV-F Analysis of Efficiency
Table IX presents a comprehensive comparison of computational efficiency and performance across different methods on the SemanticKITTI validation split. By sharing the same network architecture, our method introduces no additional runtime overhead during inference compared to UNIT, thus maintaining identical inference latency and frames per second (FPS). Notably, although 3DUIS++ achieves a competitive $\mathrm{S_{assoc}^{temp}}$ score (0.447), its computational cost is prohibitively high (latency of 74.09 s, FPS of 0.01, and training time of 189 hours), making it unsuitable for online applications. In contrast, our approach retains UNIT’s real-time capability while improving segmentation accuracy, offering a favorable balance between performance and efficiency for practical deployment.
IV-G Qualitative Results
@@
Figure 6: Qualitative examples of failure cases encountered by our method: (a) Tracking failure under fast object motion, (b) Segmentation breakdown for non-structured surfaces (wall), (c, d) A single car instance incorrectly divided into multiple IDs, highlighting limitations in point cloud representation and temporal dependency capture.
@@
Figure 5 presents qualitative comparisons on the nuScenes dataset. Our method yields more accurate and temporally consistent instance segmentation compared with pseudo-labels and UNIT. In nuScenes, UNIT introduces inconsistent predictions (highlighted with red circles), whereas our approach produces temporally consistent masks with clearer object boundaries.
While the proposed method performs well in many scenarios, we observe some limitations. For instance, when objects move at high relative speeds or are partially occluded, the temporal association mechanism struggles to maintain consistent instance IDs, leading to tracking failures, as shown in Fig.6(a). Additionally, non-structural objects like walls, fences, and vegetation, which lack clear boundaries, are sometimes over-segmented or incorrectly merged into the background, as seen in Fig.6(b). Another issue arises when sparse point clouds or insufficient temporal cues cause a single object, such as a car, to be split into multiple IDs (Fig. 6(c,d)). These challenges highlight the complexity of unsupervised instance segmentation and indicate areas for future improvements.